OpenStudy (anonymous):

construct a probability distribution for a family of 3 children. Let X represent the number of boys

OpenStudy (anonymous):

P(X = 0) = 1/8 P(X = 1) = 3/8 P(X = 2) = 3/8 P(X = 3) = 1/8

OpenStudy (anonymous):

because of the possible outcomes of having XXX then sub those as MMM and distribute if you have a female then you will add F so it will be MMF or MFM etc